OpenStudy (anonymous):

The measure of each interior angle of a regular polygon is 144°. How many sides does the polygon have? a. 9 b. 10 c. 12 d. 14

OpenStudy (anonymous):

Someone please help me?

hero (hero):

Use this formula to figure it out: \[\frac{360}{n} = 180 - A_i\] \(A_i = \) Interior Angle \(=144\)

hero (hero):

\(n = \) number of sides

OpenStudy (anonymous):

The answers 10. Right?

hero (hero):

Correct

OpenStudy (anonymous):

Thank you!